# Neural Combinatorial Optimization for Real-World Routing

### v1

@@ 6.3. Analyses
Importance of real-world data generators. We study the effect of training different models on different data generators, including the ATSP one from MatNet (Kwon et al., 2021), adding random noise to break symmetries in distance matrices, and our proposed real-world generator when testing in the real world. Table 2 demonstrates our proposed real-world data generation achieves remarkable improvements in both in-distribution and out-of-distribution settings.
@@
Table 2. Comparison of NCO solvers and their training data generators on real-world (RRNCO) data. Our RRNCO model with the proposed data generator achieves the best results.
Model Data Generator In-dist OOD (city) OOD (cluster)
@@
MatNet ATSP (MatNet) 80.86 110.70 81.04 108.30 27.78 128.23
RRNCO Noise Distance 41.35 7.72 42.01 7.98 13.66 12.20
MatNet RRNCO 39.92 3.98 40.55 4.23 12.89 5.88
RRNCO RRNCO 39.08 1.80 39.78 2.26 12.45 2.30
Ablation study. We perform an ablation study on proposed model components in Fig. 4: initial contexts with coordinates, distances, and our Contextual Gating (CG), as well as the Neural Adaptive Bias (NAB). We find CG and NAB perform the best, particularly in out-of-distribution (OOD) cases. Remarkably, in cluster distributions, the NAB shows a relative improvement greater than $15\%$.
7. Conclusion
In this paper, we introduced RRNCO to bridge the gap between synthetic and real-world vehicle routing problems. We proposed a new dataset and data generators for real-world VRPs spanning 100 diverse cities, capturing rich topological features via asymmetric distances and duration matrices with the Open Source Routing Machine. Moreover, we introduced an RL-based model employing contextual gating to integrate node and high-dimensional edge features and adaptive bias mechanisms incorporating topological features via an attention-free architecture. Our experiments demonstrate that RRNCO achieves state-of-the-art performance among NCO techniques in realistic VRP settings. By openly releasing our dataset and code, we believe RRNCO will facilitate further research toward robust, real-world routing solutions.
@@ 7. Conclusion
We make RRNCO’s code openly available at https://github.com/ai4co/real-routing-nco. The dataset is for download through the HuggingFace portal, with the link available on the Github repository. Our code is licensed under the MIT license.
Appendix A Additional Data Information
We present a comprehensive urban mobility dataset encompassing 100 cities across diverse geographical regions worldwide. For each city, we collected 1000 sampling points distributed throughout the same size urban area. The dataset includes the precise geographical coordinates (latitude and longitude) for each sampling point. Additionally, we computed and stored complete distance and travel time matrices between all pairs of points within each city, resulting in 1000×1000 matrices per city.
@@ Appendix A Additional Data Information
The cities in our dataset exhibit significant variety in their characteristics, including population size (ranging from small to large), urban layout patterns (such as grid, organic, mixed, and historical layouts), and distinct geographic features (coastal, mountain, river, valley, etc.). The dataset covers multiple regions including Asia, Oceania, Americas, Europe, and Africa. This diversity in urban environments enables comprehensive analysis of mobility patterns across different urban contexts and geographical settings.
LABEL:tab:city_details on the following page provides information about our topology dataset choices.
Appendix B Hyperparameter Details
Table 3 shows the hyperparameters we employ for RRNCO. The configuration can be changed through yaml files as outlined in RL4CO (Berto et al., 2024b), which we employ as the base framework for our codebase.
Table 3. Hyperparameters.
Hyperparameter Value
Model
Embedding dimension 128
Number of attention heads 8
Number of encoder layers 12
Normalization Instance
Use graph context False
Sample size $k$ node encoding 25
Training
Batch size 256
Train data size 100,000
Val data size 1,280
Test data size 1,280
RL algorithm REINFORCE
REINFORCE baseline POMO (Kwon et al., 2020)
Optimizer Adam
Learning rate 4e-4
Weight decay 1e-6
LR scheduler MultiStepLR
LR milestones [180, 195]
LR gamma 0.1
Max epochs 200
Data
Number of cities (training) 80
Number of cities (testing) 20
Instance size $n$ (# of locations) 100
Number of test instances 1280
Table 4. Comprehensive City Details

### v2

@@ 6.3 Analyses
We study the effect of training different models on different data generators, including the ATSP one from MatNet (Kwon et al., 2021), adding random noise to break symmetries in distance matrices, and our proposed real-world generator when testing in the real world. Table 3 demonstrates our proposed real-world data generation achieves remarkable improvements in both in-distribution and out-of-distribution settings.
@@
Table 3: Comparison of routing solvers and their training data generators on real-world data.
Method Data Gen. In-dist OOD City OOD Clust.
@@
MatNet ATSP 80.86 110.70 81.04 108.30 27.78 128.23
RRNCO Noise 41.35 7.72 42.01 7.98 13.66 12.20
MatNet Real 39.92 3.98 40.55 4.23 12.89 5.88
RRNCO Real 39.08 1.80 39.79 2.27 12.44 2.25
Generalization to stochastic VRPs
RRNCO demonstrates remarkable performance when applied to real-world topologies that surpass prior works. We further evaluate the robustness of RRNCO and generalizability to new settings, i.e., under stochastic and time-dependent traffic conditions and conduct experiments on the recently released Stochastic Multi-period Time-dependent VRP (SMTVRP) benchmark from SVRPBench (Heakl et al., 2025). We also benchmark against additional traditional baselines, including Nearest Neighbor + 2-opt (NN+2opt) (Laporte and Nobert, 1987; Potvin and Rousseau, 1995) and Ant Colony Optimization (ACO) (Dorigo et al., 2006).
As shown in Table 4, RRNCO achieves the lowest cost (601.03) while maintaining full feasibility and zero time window violations. Notably, OR-Tools fails to find feasible solutions for 75.8% of instances within the time limit, highlighting the practical advantage of neural solvers in time-constrained scenarios. These results demonstrate that RRNCO generalizes effectively to more complex, time-dependent routing problems beyond the benchmarks used for training. For more details of the dataset, baselines, and training configurations, please refer to Appendix E.
Table 4: Performance comparison on SMTVRP benchmark. Feasibility indicates the proportion of instances with valid solutions. Lower cost is better.
Method Cost Feasibility Runtime TW Violations
ACO 763.52 1.00 41.3s 0
OR-Tools 610.08 0.242 1000s 45.15
NN + 2opt 969.96 1.00 10s 0
RF 602.20 1.00 0.05s 0
GOAL 1319.00 1.00 0.28s 0
RRNCO 601.03 1.00 0.20s 0
7 Conclusion
In this paper, we introduced RRNCO, a novel Neural Combinatorial Optimization architecture bridging simplified benchmarks and real-world routing challenges. Our core contribution is a model explicitly handling asymmetric and multi-modal travel costs through two key innovations: Adaptive Node Embedding (ANE) efficiently fusing coordinates with sampled distance features, and a Neural Adaptive Bias (NAB) mechanism. This NAB represents the first approach to jointly model multiple asymmetric metrics—distance, duration, and directional angles—in real-world routing problems. To validate our model, we constructed a large-scale dataset with realistic routing data from 100 diverse cities. This dataset provides a reproducible and diverse testbed that supports future research on robust and generalizable NCO solvers. On this challenging benchmark, RRNCO achieves state-of-the-art performance among NCO methods. By releasing our model and dataset, we aim to accelerate progress in practical, deployable neural optimization solutions.
Reproducibility Statement
To facilitate reproducibility of our results, we provide complete access to our implementation, including dataset generators and training configurations. The source code is available at https://github.com/ai4co/real-routing-nco. We will additionally disclose the URL of generated data and model weights on HuggingFace. Detailed descriptions of model architectures, dataset creation, training procedures, and experimental setups are provided in the main paper and appendix.
@@ 7 Conclusion
Appendix A Detailed Model Decoder Architecture
The decoder architecture combines key elements from the ReLD and MatNet to effectively process the dense node embeddings generated by the encoder and construct solutions for vehicle routing problems. At each decoding step $t$, the decoder takes as input the row and column node embeddings ($h^{r}_{\textit{F}}$, $h^{c}_{\textit{F}}$) produced by the encoder and a context vector $h_{c}$. The composition of $h_{c}$ adapts to the problem type. For VRP variants (ACVRP and ACVRPTW), the context relies on the last visited node and the dynamic state. In contrast, for ATSP, we additionally include the embedding of the first node $a_{0}$ to explicitly anchor the tour’s origin. Formally, the context vector is defined as:
$h_{c}=\begin{cases}[h^{r}_{a_{0}},h^{r}_{a_{t-1}},D^{t}]\in\mathbb{R}^{2d_{h}+d_{\text{attr}}}&\text{for ATSP},\\ [h^{r}_{a_{t-1}},D^{t}]\in\mathbb{R}^{d_{h}+d_{\text{attr}}}&\text{for ACVRP and ACVRPTW}.\end{cases}$ (17)
where $D^{t}\in\mathbb{R}^{d_{\text{attr}}}$ represents the dynamic features derived from the state $s^{t}$. The specific contents of $D^{t}$ are tailored to the constraints of each problem variant: for ACVRP, $D^{t}$ consists of the current load; for ACVRPTW, it includes both the current load and the current time. In the case of ATSP, which is unconstrained by capacity or time windows, no dynamic features are utilized (i.e., $D^{t}=\emptyset$). To aggregate information from the node embeddings, the decoder applies a multi-head attention (MHA) mechanism, using the context vector $h_{c}$ as the query and $H^{t}\in\mathbb{R}^{|F^{t}|\times d_{h}}$ as the key and value:
$h^{\prime}_{c}=\text{MHA}(h_{c},W^{\textit{key}}h^{c}_{F},W^{\textit{val}}h^{c}_{F}).$ (18)
$h^{\prime}_{c}=h^{\prime}_{c}+\text{IDT}(h_{c}),$ (19)
$q_{c}=h^{\prime}_{c}+\text{MLP}(h^{\prime}_{c}).$ (20)
$p_{i}=\left[\text{Softmax}\left(C\cdot\tanh\left(\frac{(q_{c})^{T}W^{\ell}h^{c}_{F}}{\sqrt{d_{h}}}-\log(\text{dist}_{i})\right)\right)\right]_{i}$ (21)
Appendix B Real-World VRP Dataset Generation
Existing methodologies often require integrating massive raw datasets (e.g., traffic simulators and multi-source spatial data) – for instance, Barauskas et al. (2023) rely on simplistic synthetic benchmarks, which are either resource-intensive or lack real-world complexity (Gunawan et al., 2021).
To address these limitations, we design a three-step pipeline to create a diverse and realistic vehicle routing dataset aimed at training and testing NCO models. First, we select cities worldwide based on multi-dimensional urban descriptors (morphology, traffic flow regimes, land-use mix). Second, we develop a framework using the Open Source Routing Machine (OSRM) (Luxen and Vetter, 2011) to create city maps with topological data, generating both precise location coordinates and their corresponding distance and duration matrices between each other. Finally, we efficiently subsample these topologies to generate diverse VRP instances by adding routing-specific features such as demands and time windows, thus preserving the inherent spatial relationships while enabling the rapid generation of instances with varying operational constraints, leveraging the precomputed distance/duration matrices from the base maps. The whole pipeline is illustrated in Fig. 3.
B.1 City Map Selection
We select a list of 100 cities distributed across six continents, with 25 in Asia, 21 in Europe, 15 each in North America and South America, 14 in Africa, and 10 in Oceania. The selection emphasizes urban diversity through multiple dimensions, including population scale (50 large cities >>1M inhabitants, 30 medium cities 100K-1M, and 20 small cities <<100K), infrastructure development stages, and urban planning approaches. Cities feature various layouts, from grid-based systems like Manhattan to radial patterns like Paris and organic developments like Fez, representing different geographic and climatic contexts from coastal to mountain locations. We prioritized cities with reliable data availability while balancing between globally recognized metropolitan areas and lesser-known urban centers, providing a comprehensive foundation for evaluating vehicle routing algorithms under diverse real-world conditions. Moreover, by including cities from developing regions, we aim to advance transportation optimization research that could benefit underprivileged areas and contribute to their socioeconomic development.
B.2 Topological Data Generation Framework
Geographic boundary information
$d=2R\cdot\arcsin\left(\sqrt{\sin^{2}\left(\frac{\Delta\phi}{2}\right)+\cos(\phi_{1})\cos(\phi_{2})\sin^{2}\left(\frac{\Delta\lambda}{2}\right)}\right)$ (22)
Point sampling from road networks
Travel information computation
The travel information computation component leverages a locally hosted Open Source Routing Machine (OSRM) server (Luxen and Vetter, 2011) to calculate real travel distances and durations between sampled points, ensuring full reproducibility of results. Through the efficient get_table function in our router implementation via the OSRM table service [^3] https://project-osrm.org/docs/v5.24.0/api/#table-service, we can process a complete 1000x1000 origin-destination matrix within 18 seconds, making it highly scalable for urban-scale analyses. In contrast to commercial API-based approaches that require more than 20 seconds for 350×350 matrices (Ali and Saleem, 2024), our open-source local OSRM implementation achieves the same computations in approximately 5 seconds. Additionally, it enables the rapid generation of multiple instances from small datasets with negligible computational cost per iteration epoch. The RRNCO framework finally processes this routing data through a normalization strategy that addresses both unreachable destinations and abnormal travel times. This step captures real-world routing complexities, including one-way streets, turn restrictions, and varying road conditions, resulting in asymmetric distance and duration matrices that reflect actual urban travel patterns. All computations are performed locally [^4] Our framework can also be extended to include real-time commercial map API integrations and powerful traffic forecasting to obtain better-informed routing (Wang and Xu, 2011; Qu and Wu, 2025), which we leave as future works., allowing for consistent results and independent verification of the analysis pipeline.
B.3 VRP Instance Subsampling
From the large-scale city base maps, we generate diverse VRP instances by subsampling a set of locations along with their corresponding distance and duration matrices, allowing us to generate an effectively unlimited number of instances while preserving the underlying structure. The subsampling process follows another three-step procedure:
1. Index Selection: Given a city dataset containing $N_{\text{tot}}$ locations, we define a subset size $N_{\text{sub}}$ representing the number of locations to be sampled for the VRP instance. We generate an index vector $\mathbf{s}=(s_{1},s_{2},\dots,s_{N_{\text{sub}}})$ where each $s_{i}$ is drawn from $\{1,\dots,N_{\text{tot}}\}$, ensuring unique selections. 2. Matrix Subsampling: Using $\mathbf{s}$, we extract submatrices from the precomputed distance matrix $D\in\mathbb{R}^{N_{\text{tot}}\times N_{\text{tot}}}$ and duration matrix $T\in\mathbb{R}^{N_{\text{tot}}\times N_{\text{tot}}}$, forming instance-specific matrices $D_{\text{sub}}=D[\mathbf{s},\mathbf{s}]\in\mathbb{R}^{N_{\text{sub}}\times N_{\text{sub}}}$ and $T_{\text{sub}}=T[\mathbf{s},\mathbf{s}]\in\mathbb{R}^{N_{\text{sub}}\times N_{\text{sub}}}$, preserving spatial relationships among selected locations. 3. Feature Generation: Each VRP can have different features. For example, in Asymmetric Capacitated VRP (ACVRP) we can generate a demand vector $\mathbf{d}\in\mathbb{R}^{N_{\text{sub}}\times 1}$, such that $\mathbf{d}=(d_{1},d_{2},\dots,d_{N_{\text{sub}}})^{\top}$, where each $d_{i}$ represents the demand at location $s_{i}$. Similarly, we can extend to ACVRPTW (time windows) represented as $\mathbf{W}\in\mathbb{R}^{N_{\text{sub}}\times 2}$, where $\mathbf{W}=\{(w_{1}^{\text{start}},w_{1}^{\text{end}}),\dots,(w_{N_{\text{sub}}}^{\text{start}},w_{N_{\text{sub}}}^{\text{end}})\}$, defining the valid service interval for each node.
Unlike previous methods that generate static datasets offline (Duan et al., 2020; Ali and Saleem, 2024), our RRNCO generation framework dynamically generates instances on the fly in few milliseconds, reducing disk memory consumption while maintaining high diversity. Fig. 3 illustrates the overall process, showing how a city map is subsampled using an index vector to create VRP instances with distance and duration matrices enriched with node-specific features such as demands and time windows. Our approach allows us to generate a (arbitrarily) large number of problem instances from a relatively small set of base topology maps totaling around 1.51.5GB, in contrast to previous works that required hundreds of gigabytes of data to produce just a few thousand instances.
B.4 Additional Data Information
@@
We present a comprehensive urban mobility dataset encompassing 100 cities across diverse geographical regions worldwide. For each city, we collected 1000 sampling points distributed throughout the same size urban area. The dataset includes the precise geographical coordinates (latitude and longitude) for each sampling point. Additionally, we computed and stored complete distance and travel time matrices between all pairs of points within each city, resulting in 1000×1000 matrices per city. The cities in our dataset exhibit significant variety in their characteristics, including population size (ranging from small to large), urban layout patterns (such as grid, organic, mixed, and historical layouts), and distinct geographic features (coastal, mountain, river, valley, etc.). The dataset covers multiple regions including Asia, Oceania, Americas, Europe, and Africa. This diversity in urban environments enables comprehensive analysis of mobility patterns across different urban contexts and geographical settings. LABEL:tab:city_details on the following page provides information about our topology dataset choices.
Table 5: Comprehensive City Details
@@
Appendix C Additional Experimental Details
C.1 Hyperparameter Details
Table 6 shows the hyperparameters we employ for RRNCO. The configuration can be changed through yaml files as outlined in RL4CO (Berto et al., 2025a), which we employ as the base framework for our codebase.
Table 6: Hyperparameters for RRNCO.
Hyperparameter Value
Optimizer Adam
Learning Rate $4\times 10^{-4}$
LR Decay Schedule 0.1 at epochs 180, 195
Batch Size 256
Instances per Epoch 100,000
Embedding Dimension 128
Feedforward Dimension 512
AAFM Layers 12
Clipping $C$ 10
C.2 Testing Dataset
For each dataset of the main experiments Section 6.2, we use subsampling as described in Section B.3 and sample 1,280 instances for each test set. Whole test sets and seeds are provided in the shared code for reproducibility.
C.3 Baselines Details
All evaluations are conducted on an NVIDIA A6000 GPU paired with an Intel(R) Xeon(R) CPU @ 2.20GHz. For neural methods, we employ the provided models and code in the original repositories to ensure fairness and reproducibilty.
We evaluate all traditional solvers on a single CPU core sequentially. To reflect a realistic deployment scenario, we do not perform instance-specific hyperparameter tuning, instead relying on the robust default configurations provided by each library. For LKH-3 (Helsgaun, 2017), we utilize the standard parameter set from the official distribution (e.g., PATCHING_C=3, PATCHING_A=2 for VRPs), imposing a TIME_LIMIT of 5 seconds per instance. For PyVRP (Wouda et al., 2024), we employ the default Hybrid Genetic Search (HGS) parameters specifically nb_elite=4 and generation_size=40 with a 20-second time limit via MaxRuntime. Similarly, for Google OR-Tools (Perron and Didier, 2024), we configure the routing model to use the GUIDED_LOCAL_SEARCH metaheuristic initialized with PATH_CHEAPEST_ARC, restricted to a 20-second budget. This protocol, consistent with Liu et al. (2024), results in a total evaluation time of approximately 7 hours for the test set (1,280 instances).
Appendix D Use of Large Language Models
Large language models were employed solely as general-purpose writing assistants. Their use was restricted to refining phrasing, improving clarity, and correcting grammar in draft versions of the manuscript. All research ideas, methodologies, analyses, results, and interpretations were conceived, executed, and validated exclusively by the authors. Any text generated with the assistance of LLMs was thoroughly reviewed, edited, and integrated by the authors to ensure accuracy, correctness, and compliance with academic standards.
Appendix E Additional Material on Stochastic Benchmarks
We further evaluate RRNCO on routing scenarios with time-varying travel conditions. Specifically, we benchmark on the Stochastic Multi-period Time-dependent VRP (SMTVRP), which incorporates realistic traffic dynamics following the simulation protocol of SVRPBench (Heakl et al., 2025).
E.1 Benchmark Setup
Dynamic Travel Time Model
Real-world travel times exhibit significant temporal variation due to congestion patterns, stochastic delays, and unexpected incidents. We model the travel time from node $a$ to $b$ departing at time $t$ as:
$T(a,b,t)=T_{\text{base}}(a,b)+T_{\text{congestion}}(a,b,t)+T_{\text{incident}}(t),$ (23)
where $T_{\text{base}}(a,b)=D(a,b)/V$ represents the free-flow travel time based on Euclidean distance $D(a,b)$ and average speed $V$. The congestion component captures systematic daily patterns:
$T_{\text{congestion}}(a,b,t)=B(a,b,t)\cdot R(t),$ (24)
with $B(a,b,t)$ encoding deterministic congestion and $R(t)$ introducing stochastic variability.
The congestion factor combines temporal and spatial dependencies:
$B(a,b,t)=\alpha\cdot\underbrace{\left[\beta+\gamma\sum_{p\in\{\text{am},\text{pm}\}}\mathcal{G}(t;\mu_{p},\sigma_{p})\right]}_{F_{\text{time}}(t)}\cdot\underbrace{\left(1-e^{-D(a,b)/\lambda}\right)}_{F_{\text{dist}}(D)},$ (25)
where $\mathcal{G}(t;\mu,\sigma)$ denotes a Gaussian kernel. The bimodal temporal structure with peaks at $\mu_{\text{am}}=8$ and $\mu_{\text{pm}}=17$ reflects typical morning and evening rush hours. The distance factor captures the empirical observation that longer trips have higher congestion exposure.
To model traffic variability, $R(t)$ follows a log-normal distribution with time-dependent parameters:
$R(t)\sim\text{LogNormal}\big(\mu_{R}(t),\sigma_{R}(t)\big),$ (26)
where both $\mu_{R}(t)$ and $\sigma_{R}(t)$ increase during peak hours to reflect heightened uncertainty.
Incident-induced delays are modeled as:
$T_{\text{incident}}(t)=\mathbb{1}[\text{incident at }t]\cdot\Delta_{\text{incident}},$ (27)
where incident occurrences follow a time-inhomogeneous Poisson process with elevated rates during nighttime hours ($\mu_{\text{night}}=21$), and delay durations $\Delta_{\text{incident}}\sim U(0.5,2.0)$ hours align with industry clearance statistics.
Time Window Generation
Customer availability windows are sampled to reflect realistic delivery scenarios. We distinguish between two customer types with distinct temporal preferences:
$T_{\text{start}}^{\text{(res)}}\sim 0.5\cdot\mathcal{N}(\mu_{\text{am}},\sigma_{\text{am}}^{2})+0.5\cdot\mathcal{N}(\mu_{\text{pm}},\sigma_{\text{pm}}^{2}),$ (28)
for residential customers with morning ($\mu_{\text{am}}=480$ min) and evening ($\mu_{\text{pm}}=1140$ min) availability peaks, and
$T_{\text{start}}^{\text{(com)}}\sim\mathcal{N}(\mu_{\text{biz}},\sigma_{\text{biz}}^{2}),$ (29)
for commercial customers centered on business hours ($\mu_{\text{biz}}=780$ min). Window durations are sampled uniformly within problem-specific bounds. Complete parameter specifications follow SVRPBench (Heakl et al., 2025).
E.2 Baseline Configurations
We evaluate a set of baseline methods on the stochastic multi-period time-dependent VRP (SMTVRP) benchmark to ensure rigorous and fair comparisons. All methods are configured following standard practices in SVRPBench and prior VRP literature, with adjustments only when necessary to accommodate time-dependent travel information.
ACO (Ant Colony Optimization)
We adopt a conventional parameterization commonly used in the SVRP literature. To balance solution quality and computational efficiency in dynamic environments, we set the number of ants to 50 and the maximum number of iterations to 100. The pheromone-related parameters are fixed to $\alpha=1.0$, $\beta=2.0$, and evaporation rate $\rho=0.5$. At each temporal update in SMTVRP, heuristic costs are recomputed while pheromone trails are preserved to maintain stability across time periods.
OR-Tools
We use the guided local search (GLS) implementation included in SVRPBench with a strict time limit of 1000 seconds per instance. This extended budget allows the solver to explore large neighborhoods under dynamic updates. Following each change in the time-dependent duration matrix, OR-Tools is restarted from the most recent feasible solution when possible, but no custom tuning beyond the default GLS configuration is introduced.
NN + 2-opt
In accordance with the SVRPBench protocol, a Nearest Neighbor (NN) heuristic is first used to construct an initial solution, followed by a 2-opt local improvement phase. The search terminates upon convergence or after reaching a maximum runtime of 10 seconds per instance. Under dynamic conditions, NN + 2-opt is rerun at every update step to preserve consistency with the updated travel-time information.
GOAL
We evaluate GOAL using its official pre-trained checkpoint trained on CVRPTW instances. No additional training or adaptation is performed. For each temporal update, GOAL re-evaluates its autoregressive decoding process under the current duration matrix, thereby testing its zero-shot generalization capacity to dynamic inputs.
RouteFinder
We initialize RouteFinder from its official pre-trained model and apply Efficient Active Learning (EAL) for 20 epochs to adapt the policy to the SMTVRP distributions. After EAL adaptation, inference proceeds in a closed-loop manner: at each step, the updated travel-time matrix is processed, and actions are generated based on the current state of the dynamic environment.
E.3 Multi-Snapshot NAB for Stochastic Routing
To extend RRNCO to robustly handle time-dependent and stochastic traffic variations, we introduce a Multi-Snapshot variant of the Neural Adaptive Bias (NAB). While the standard NAB fuses a single duration matrix, real-world traffic is highly dynamic and uncertain. To address this, we integrate a dynamics-informed traffic simulation directly into the encoder pipeline and extend the NAB architecture to process multiple temporal ”snapshots” of traffic conditions simultaneously.
E.3.1 Dynamics-Informed Input Generation
To approximate the stochastic nature of real-world travel times, we employ the dynamics-informed traffic modeling framework proposed in SVRPBench (Heakl et al., 2025). Instead of a single static duration matrix, the model takes as input a set of $K$ distinct duration snapshots $\{\mathbf{T}_{1},\dots,\mathbf{T}_{K}\}$. In our experiments, we set $K=3$ to represent distinct traffic regimes (e.g., morning peak, noon, evening peak).
These snapshots are generated using Gaussian kernels centered at learnable “time anchors” to model congestion peaks, combined with stochastic noise injection. Specifically, for a snapshot $k$, the duration $t_{ij}^{(k)}$ is derived by modulating the base travel time with a congestion factor drawn from a time-dependent Gaussian distribution, multiplicative noise sampled from a LogNormal distribution to simulate flow variability, and sparse additive noise modeled via a Poisson process to account for random incidents (accidents). This ensures the model receives a global, “look-ahead” view of potential traffic states and their associated uncertainties.
E.3.2 Multi-Snapshot Gating Mechanism
We extend the Neural Adaptive Bias (NAB) module to aggregate these heterogeneous temporal views. The Multi-Snapshot NAB employs a unified gating network that dynamically weighs the importance of the static spatial structure against the variable traffic conditions.
First, we project the static Distance matrix $\mathbf{D}$, the Relative Angle matrix $\mathbf{\Phi}$, and each of the $K$ dynamic Duration snapshots into a shared high-dimensional embedding space:
$\mathbf{D}_{\text{emb}}$ $=\text{ReLU}(\mathbf{D}\mathbf{W}_{D})\mathbf{W}^{\prime}_{D},$ (30)
$\mathbf{\Phi}_{\text{emb}}$ $=\text{ReLU}(\mathbf{\Phi}\mathbf{W}_{\Phi})\mathbf{W}^{\prime}_{\Phi},$ (31)
$\mathbf{T}^{(k)}_{\text{emb}}$ $=\text{ReLU}(\mathbf{T}_{k}\mathbf{W}_{T})\mathbf{W}^{\prime}_{T},\quad\forall k\in\{1,\dots,K\}.$ (32)
To fuse these $K+2$ feature maps, we compute a set of scalar importance weights via a learned gating network. Let $\mathbf{E}_{\text{concat}}=[\mathbf{D}_{\text{emb}};\mathbf{\Phi}_{\text{emb}};\mathbf{T}^{(1)}_{\text{emb}};\dots;\mathbf{T}^{(K)}_{\text{emb}}$] denote the concatenation of all embeddings. We compute the gating weights $\mathbf{g}\in\mathbb{R}^{2+K}$ via a softmax function:
$\mathbf{g}=\text{softmax}\left(\frac{\mathbf{E}_{\text{concat}}\mathbf{W}_{G}}{\exp(\tau)}\right),$ (33)
where $\mathbf{W}_{G}$ is a learnable weight matrix and $\tau$ is the temperature parameter. The final fused representation $\mathbf{H}_{\text{fused}}$ is computed as a probability-weighted sum of the static and dynamic features:
$\mathbf{H}_{\text{fused}}=g_{1}\odot\mathbf{D}_{\text{emb}}+g_{2}\odot\mathbf{\Phi}_{\text{emb}}+\sum_{k=1}^{K}g_{k+2}\odot\mathbf{T}^{(k)}_{\text{emb}}$. (34)
Similar to the standard NAB, this fused representation is projected to a scalar bias matrix $\mathbf{A}=\mathbf{H}_{\text{fused}}\mathbf{w}_{\text{out}}$. This bias $\mathbf{A}$ modulates the attention scores in the Adaptation Attention Free Module (AAFM), guiding the solver to avoid routes that are consistently congested across the simulated snapshots or to leverage time windows where traffic is predicted to be lighter.
E.4 Training Details
RRNCO (NAB)
The standard RRNCO model equipped with the Neural Adaptive Bias (NAB) is trained for 100 epochs on static instances sampled from the real-world generator introduced in the main paper. Each epoch consists of 100,000 training instances, with a batch size of 256 and the Adam optimizer with a learning rate of $4\times 10^{-4}$. The learning rate is decayed following the schedule described in Table 6 of the appendix. No fine-tuning is performed on the SMTVRP benchmark; instead, the model is directly evaluated in a dynamic closed-loop setting.
RRNCO (Multi-Snapshot NAB)
As detailed in Section E.3, this variant integrates temporal variability by augmenting the Neural Adaptive Bias (NAB) module with a dynamics-informed traffic simulator and a temporal gating mechanism designed to statically fuse multiple duration snapshots. In our experiments, the model receives $K=3$ temporal duration snapshots at training time, corresponding to short-horizon predictions of traffic fluctuations. The model is trained for 100 epochs under the same reinforcement learning setup as the base RRNCO. During dynamic evaluation on SMTVRP, the NAB module processes and fuses the $K$ snapshots once during the initial encoding step. This process enables the resultant node representations to embed a global context of potential traffic variations, allowing the autoregressive decoder to anticipate near-future congestion patterns without requiring dynamic re-encoding at every decision step.
Inference Protocol
Both RRNCO variants operate autoregressively in a closed-loop dynamic environment. After each routing action, the updated time-dependent duration matrix is provided by the SMTVRP simulator. The model recomputes its route continuation based on this updated information. A decoding augmentation factor of 8 is applied during inference to reduce variance and enhance route stability across dynamic updates.
E.5 Results for Multi-Snapshot NAB
Table 7 demonstrates that RRNCO with Multi-Snapshot NAB achieves state-of-the-art performance with a cost of 594.19, surpassing both the standard NAB (601.03) and RouteFinder (602.20). This improvement confirms that fusing multiple temporal snapshots allows the model to better anticipate traffic stochasticity. On the other hand, OR-Tools fails to find feasible solutions for 75.8% of instances despite a 1000s budget. RRNCO maintains 100% feasibility and zero time window violations, offering superior robustness and speed (0.20s) with virtually no additional inference latency compared to the standard model.
Table 7: Performance comparison on SMTVRP benchmark. Feasibility indicates the proportion of instances with valid solutions. Lower cost is better.
Method Cost Feasibility Runtime TW Violations
ACO 763.52 1.00 41.3s 0
OR-Tools 610.08 0.242 1000s 45.15
NN + 2opt 969.96 1.00 10s 0
RF 602.20 1.00 0.05s 0
GOAL 1319.00 1.00 0.28s 0
RRNCO (NAB) 601.03 1.00 0.20s 0
RRNCO (Multi-Snapshot NAB) 594.19 1.00 0.20s 0